OpenStudy (anonymous):

What is the slope-intercept form of the line that passes through the points (-4,-2) and (-2, 6)?

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (gorv):

slope = (y2-y1)/(x2-x1) =(6-(-2))/(-2-(-4))=8/2=4 slope=4 take any point out of these two y-6=4(x-(-2)) y-6=4x+8 y=4x+14 this is slope intercept form slope =4 y intercept =14 @hwilkx3